OpenStudy (anonymous):

The percentage of students graduating from high school in an urban area in Maryland decreased from 87.4% in 2005 to 81.8% in 2009. Predict the percentage of students who will graduate from high school in 2012

OpenStudy (anonymous):

Annual rate = (87.4 - 81.8)/4 = 1.4 Rate in 2012 will be 81.8 - (1.4 * 3) = 77.6%

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

No problem buddy!!!you rocks..!

OpenStudy (anonymous):

thanks!